OpenStudy (anonymous):

Find all solutions to sin^2-cos^2=1

OpenStudy (anonymous):

sin^2-cos^2=1 sin^2-cos^2=sin^2x+cos^2x -cos^2=+cos^2x Can you take the problem from here ?

OpenStudy (anonymous):

Possibly is there a next step or is the only solution just 0?

OpenStudy (anonymous):

idk how to go on... sorry -:(

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

thank you for the start!!

OpenStudy (loser66):

-cos^2 = cos^2 2cos^2 =0 iff cos^2 =0 or cos=0 when x = 0 + 2kpi

OpenStudy (loser66):

clear?

OpenStudy (anonymous):

ohhh alright that makes sense. thank you so much!

OpenStudy (loser66):

oh, sorry when x = pi/2 +kpi, not 0, I think of sin

OpenStudy (anonymous):

ohh alright. thanks again